# HARMONY: Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models

### v1

@@ 5.1 Experimental Setup
Datasets and Models We evaluate our work on three VQA datasets, A-OKVQA schwenk2022okvqa , VizWiz gurari2018vizwiz and a PathVQA he2020pathvqa . The A-OKVQA datasets require reasoning and common sense alongside visual information. The VizWiz dataset covers a challenging setup, where each image is taken by a blind/visually impaired individual and accompanied by spoken questions about the images. These questions are then transcribed. PathVQA is a medical imaging VQA dataset. To train the scoring function, we leverage the training splits of the corresponding datasets. We use a 80% and 20% split to construct a train and a validation split, respectively. To evaluate the performance of the scoring function, we use the validation split given with the dataset as a test split. We provide further details of datasets and training strategy/hyper-parameters in Appendix B.2. We evaluate our method on three open-sourced VLMs: LLaVA-7b liu2023visual , LLaVA-13b liu2024improved and InstructBLIP huang2023visual . InstructBLIP uses FlanT5-XL raffel2020exploring as the LLM backbone. All these models are instruction-tuned on the VQA task.
@@
Baselines We include a range of black-box approaches, including length-normalized confidence malinin2020uncertainty , first-token confidence zhao2024first , and self-eval srinivasan2024selective , Entropy, Semantic Entropy, and Cluster Entropy farquhar2024detecting . Semantic Entropy, which measures consistency among semantically similar answers and can be regarded as an alternative implementation of khan2024consistency . Finally, we consider supervised training-based methods, including MSF (Multimodal Selection Function) whitehead2022reliable , which trains an MLP on hidden representations of the prompt and generated answer, and LARS yaldiz2024not , which trains a transformer architecture on the token probabilities predicted by the base model.
@@ 5.2 Results
Table 1: AUROC and PRR scores on A-OKVQA and VizWiz dataset
LLaVA - 7b LLaVA - 13b Instruct-BLIP
UE Method AUROC(%) PRR(%) AUROC(%) PRR(%) AUROC(%) PRR(%)
@@
Length-Normalized Confidence 74.55 61.45 77.50 67.04 74.13 56.60
First Token Confidence 69.39 33.16 72.96 41.35 75.09 58.47
Self-Eval Confidence 71.53 54.48 63.04 54.43 76.12 62.75
Entropy 61.38 35.65 67.57 49.23 54.15 30.15
Semantic Entropy 78.39 68.48 80.83 69.89 73.72 52.20
Cluster Entropy 69.87 52.27 68.90 50.89 71.00 51.11
MSF 78.66 67.01 77.64 67.07 79.93 67.31
LARS 79.90 68.95 81.46 73.70 80.07 68.31
HARMONY [Ours] 83.99 75.05 83.72 77.09 81.73 72.03
(+4.09) (+6.10) (+2.26) (+3.36) (+1.66) (+3.72)
@@
Length-Normalized Confidence 71.94 44.30 75.61 55.32 77.57 75.51
First Token Confidence 69.76 43.06 71.52 46.65 76.00 56.48
Self-Eval Confidence 63.09 30.62 67.22 42.06 73.43 52.01
Entropy 33.84 15.16 41.54 19.32 38.31 28.72
Semantic Entropy 64.04 21.82 70.89 38.21 66.43 27.90
Cluster Entropy 59.86 18.83 65.23 30.25 66.09 30.40
MSF 85.66 74.43 86.43 74.42 86.62 73.67
LARS 80.50 64.13 85.29 73.09 86.34 72.14
HARMONY [Ours] 87.26 76.83 88.71 79.68 86.63 73.75
(+1.60) (+2.40) (+2.28) (+5.26) (+0.01) (+0.08)
5.2.1 UE Performance
@@
We present the results of comparison of our method HARMONY with other UE baselines on A-OKVQA and VizWiz datasets in Table 1. Among the black-box methods, for the AOKVQA daatset, we observe that SE is a strong baseline among all the black-box methods considered in our study for LLaVa model series. However, for InstructBLIP model, self-eval performs relatively better. Note that self-eval requires two forward passes, whereas SE requires five forward passes. Therefore, they can be expensive, requiring multiple forward passes of the 7B or 13B models. On the contrary, our method, HARMONY, uses VisualBERT, consisting of 113M parameters, and requires only one forward pass to yield the reliability score. Further, for VizWiz dataset consisting of unanswerable question, LCS performs better across all models. Among trainable functions, multimodal selection function (MSF) and LARS improves upon the LNC consistently across all datasets, and all models. However, our proposed method HARMONY consistently performs on par or better than LARS and MSF achieving upto 4% increase in AUROC scores and 6% increase in the PRR scores.
Medical imaging Dataset: We also report the UE performance on medical imaging dataset, PathVQA dataset. For this dataset, we find that the output probability based scoring functions, such as LNC, first token, self-eval, entropy, semantic entropy, and cluster entropy perform significantly lower than the learnable functions as shown in 2. Further, HARMONY achieves state-of-the-art performance indicating its ability to capture visual-language uncertainty in medical domain as well.
@@ 5.2.2 Selective Prediction Performance
Here, we present the evaluation of the learnable scoring UE baselines on the selective prediction task. Before we compare these methods, it is important to mention that for trainable functions, we select the best model checkpoints based on the AUROC scores. However, a user may choose a different criterion such as ER to achieve higher performance on this downstream task. The objective here is to present a practical use case, and compare the performance of the learnable UE methods.
LLaVa - 7b LLaVa - 13b Instruct-BLIP
UE Method ER(%) C@R=10% C@R=20% ER(%) C@R=10% C@R=20% ER(%) C@R=10% C@R=20%
AOKVQA
MSF 49.17 43.75 80.17 53.19 50.56 89.52 38.15 32.66 60.08
LARS 49.78 50.65 82.53 53.71 61.83 90.48 37.73 31.27 61.83
HARMONY [Ours] 52.31 60.61 85.59 55.90 64.80 92.23 38.25 36.77 63.32
VizWiz
MSF 21.30 15.00 34.15 23.47 11.90 37.76 13.01 9.15 19.84
LARS 15.72 4.91 21.53 19.14 8.86 29.08 12.82 6.85 17.13
HARMONY [Ours] 21.93 16.37 36.03 24.69 20.24 41.11 13.38 9.23 19.94
First, we report how many questions are covered by our method at 10% and 20% risk levels. The larger the number, the better performance. We observe that our proposed method consistently achieves higher coverage across various models, and datasets. We also report ER metric, which represents a better tradeoff between coverage and risk due to a penalty on the incorrectly covered question. For the comparison, we select a threshold for each method giving best ER on the validation split of calibration set, and compute effective reliability using that threshold on the test set. For this metric, our method either performs similar or outperforms other methods achieving up to 2.5% higher score. This highlights its potential to yield higher coverage while inuring lower risks. As an example, we present some sample questions in 6 and compare the decision predictions on the trainable functions. While training on either output distributions or hidden representations alone can lead to contradictory or consistently incorrect decisions, leveraging both simultaneously results in better decision functions.
@@
5.3 Out-of-Distribution Generalization Performance
To evaluateout-of-distribution (OOD) generalization, we test the LLaVA-13B model trained on the A-OKVQA dataset, which focuses on visual reasoning, by introducing OOD samples from the OKVQA dataset. The experimental setup, illustrated in Figure 3, progressively increases the proportion of OOD samples in the evaluation set. Specifically, the x-axis denotes the percentage of OOD data, where, for example, 33% corresponds to a test mixture containing 33% OKVQA (OOD) samples and 66% A-OKVQA (in-distribution) samples. As the proportion of OOD data increases, we observe a consistent decline in AUROC scores across all supervised baselines, including model confidence based scoring such as Length Normalized Confidence indicating a degradation in uncertainty calibration under distributional shift. In contrast, HARMONY consistently achieves higher AUROC values, demonstrating relatively better generalization to unseen data distributions.
@@ 5.3.1 Ablation across model layers and input signals
Our hypothesis is that internal layers have a signal of multimodal reliability, however, it is not clear which layer would provide the best signal. Previous works have highlighted that inner layers (layers closer to the input) focus more on extracting lower-level information from the input, while outer layers (layers closer to the model output) are mostly focused on the next token generation azaria2023internal . Therefore, for all the models, we ablate over every fourth layer for both MSF, and our method. We report the best performing layer results in Table 1. We observe that for LLaVa-7b and 13b models, inner layers (layer 16 and layer 22) yield the best AUROC performance for our method across both datasets. Further, for InstructBLIP, we find the outer-most layer performs the best.
@@
We conduct an ablation study on partial input signals, generated tokens (Text), token-level probabilities (Prob), and hidden states (HS), using the LLaVA-13B model on the A-OKVQA dataset. We use VisualBERT as the scoring function transformer architecture for this study. As shown in 5, we find that text (question and generated answer) without the token probabilities and hidden states yield no significant UE estimate. However, addition of token probabilities accompanying each token text helps the scoring function learn a better UE estimate. Likewise, if we use text and hidden states of the generated tokens as input signals, they yield results comparable to the text and probability baseline. However, combining all three sources of information as proposed in our workyield significantly better reliability estimates.
Figure 4: Effect of calibration data size on the UE performance [A-OKVQA dataset].
5.4 Effect of Calibration Data
For learnable functions, it important to conduct an ablation study to investigate how the function’s performance scales with the amount of calibration data. For this, we vary the calibration set size with LLaVA-13b model and report AUROC. As shown in Figure 4, increasing the calibration data size consistently enhances the AUROC scores across all learnable baselines. We also observe that HARMONY requires at least 2,000 samples to achieve performance comparable to LNC, and approximately 6,000 samples to surpass existing SOTA baselines. This trend highlights that while learnable approaches benefit significantly from more calibration data, their relative advantage becomes more apparent only once sufficient data are available to capture the diversity of uncertainty patterns in multimodal settings.
@@ Conclusion
This work introduces a novel uncertainty estimation method HARMONY for Vision-Language Models that effectively combines hidden activation representations with output token probabilities and generated token text. By jointly leveraging model internal states, generated tokens and output beliefs in a sequential fashion, our proposed framework provides a more holistic reliability assessment, complementing probability-based and representation-based approaches. Our extensive experiments on AOKVQA and VizWiz datasets demonstrate that our method significantly improves UE for multimodality, achieving up to 4% AUROC, 6% PRR, and 2.5% effective reliability score improvements over existing state-of-the-art UE methods.
Appendix A Motivation
This section further elaborates on the UE challenges presented in section 4.
(a) Token-Level Semantics
(b) Unanswerable vs. Answerable
(c) Average prob. and accuracy gap
Figure 5: Vision-language models often struggle to produce reliable confidence estimates for their generations. Figure 5(a) illustrates that certain tokens carry greater visual significance (for instance, the token “up” in this example). Figure 5(b) shows that the confidence distributions for unanswerable, visually uncertain samples can be skewed toward higher confidence regions ($\geq$ 40%). Finally, Figure 5(c) compares the average predicted probabilities and accuracies for visually uncertain questions, highlighting a notable mismatch between the two.
A.1 Token-Level Semantic and Visual Importance
Identifying an effective aggregation strategy through heuristics is challenging problem due to inter-token dependencies and various inherent biases for the given vision and text context. This can be illustrated by the example presented in 5(a). In response to the question ‘Where is the man looking?’ for the given image, the model generates the answer ‘The man is looking up’. In this case, the token up holds more semantic significance than the other tokens such as ’the’, ’is’, and ’looking’. Further, its relevant uncertainty score should weigh more in the aggregation formulation. Hence, identifying these inter-token dependencies and the underlying token-level semantic and visual significance on-the-fly is a challenging problem.
A.2 Biased Probabilities and Language Priors
To show this issue of biased probability and Language priors in VLMs, we perform visual question answering inference on the VizWiz dataset, a visual question answering (VQA) benchmark comprising images collected by blind people with their spoken questions transcribed by annotators. Due to the nature of the dataset, it includes a significant portion of unanswerable questions. For such questions, we expect model response to be on the low-confidence density region to reflect visual uncertainty. However, as shown in Figure 5(b), we observe a large density of length-normalized confidence scores to be above 40% confidence region, depicting the issue of biased probabilities, and tendency of model to hallucinate wrong answers with high confidence on visually uncertain question-image pairs. We also record model’s average accuracy versus average probability scores on unanswerable and answerable questions for LLaVa-1.5-7b model, and find that for visually uncertain answers ( unanswerable questions) the gap between average probability score and average accuracy is significantly high, >30%. This highlights the issue of model hallucinating wrong answers with high confidence under visual uncertainty. Given the observation that output probabilities can be biased, it is necessary to collect more signals of reliability from the model, reflecting visual grounding.
A.3 Internal Hidden States as a Visual Understanding Signal
Model internal hidden states have been studied extensively to predict truthfulness [4], fairness [35], and sometimes control/steer the model response towards certain concepts such as safety [36]. They have also been studied to interpret multimodality misalignment [1]. Various works suggest that model’s attention to its visual and text tokens varies across layers. Internal or middle layers attend highest to the vision input, whereas output layers pay more attention to input text query, collect their thoughts and decide what to say/generate [15, 1]. To design supervised scoring functions, they have been leveraged to predict reliability on MLP-based functions for LLM [4] and VLMs [31]. Where these functions show great effectiveness as predictors, they overlook the query text, generated response, and the confidence scores model assigns to these generated tokens, which also contain information of vision-text alignment, and reliability.
Given these observations, we hypothesize that the use of both hidden state representations and explicit token-level uncertainty provides a more holistic measure of reliability. Internal activations capture latent uncertainty, revealing model’s understanding of the vision context and how it aligns visual and textual information, while output probabilities track confidence shifts throughout generation, providing deeper insights into model’s uncertainty and trustworthiness.
Appendix B Additional Experiments and Details
B.1 Dataset Information
A-OKVQA comprises 17.1K train and 1.1K validation samples. We use the train split as a calibration dataset. For training our scoring function, we further divide the calibration dataset between 80% train and 20% validation data partition. To evaluate our method, we use the 1.1K validation split provided with the original dataset as test set. We use similar calibration data split setup for VizWiz and PathVQA datasets. Further, for inference, for all the models, we use a prompt of <image>question, please provide a single word or short sentence answer. ASSITANT:.
B.2 Training Strategy

### v2

@@ 5.1 Experimental Setup
Datasets and Models We evaluate our work on two VQA datasets, A-OKVQA schwenk2022okvqa and VizWiz gurari2018vizwiz. The A-OKVQA datasets require reasoning and common sense alongside visual information. The VizWiz dataset covers a challenging setup, where each image is taken by a blind/visually impaired individual and accompanied by spoken questions about the images. These questions are then transcribed. We also include experiments on PathVQA he2020pathvqa which is a medical imaging VQA dataset. To train the scoring function, we leverage the training splits of the corresponding datasets. We use a 80% and 20% split to construct a train and a validation split, respectively. To evaluate the performance of the scoring function, we use the validation split given with the dataset as a test split. We provide further details of datasets and training strategy/hyper-parameters in Appendix A.2. We evaluate our method on four open-sourced VLMs: LLaVA-7b liu2023visual, LLaVA-13b liu2024improved, InstructBLIP huang2023visual and Qwen-VL bai2023qwen. InstructBLIP uses FlanT5-XL raffel2020exploring as the LLM backbone. All these models are instruction-tuned on the VQA task.
Table 1: AUROC and PRR scores on A-OKVQA and VizWiz dataset
LLaVA - 7b LLaVA - 13b Instruct-BLIP Qwen-VL
UE Method AUROC(%) PRR(%) AUROC(%) PRR(%) AUROC(%) PRR(%) AUROC(%) PRR(%)
AOKVQA
Length-Normalized Confidence 74.55 61.45 77.50 67.04 74.13 56.60 65.53 33.83
First Token Confidence 69.39 33.16 72.96 41.35 75.09 58.47 63.59 28.28
Self-Eval Confidence 71.53 54.48 63.04 54.43 76.12 62.75 58.06 10.81
Entropy 61.38 35.65 67.57 49.23 54.15 30.15 69.11 48.48
Semantic Entropy 78.39 68.48 80.83 69.89 73.72 52.20 59.05 11.05
Cluster Entropy 69.87 52.27 68.90 50.89 71.00 51.11 59.05 18.08
MSF 78.66 67.01 77.64 67.07 79.93 67.31 76.82 64.94
LARS 79.90 68.95 81.46 73.70 80.07 68.31 74.63 54.37
HARMONY [Ours] 83.99 75.05 83.72 77.09 81.73 72.03 81.84 74.10
(+4.09) (+6.10) (+2.26) (+3.36) (+1.66) (+3.72) (+5.02) (+9.14)
VizWiz
Length-Normalized Confidence 71.94 44.30 75.61 55.32 77.57 75.51 71.97 45.10
First Token Confidence 69.76 43.06 71.52 46.65 76.00 56.48 70.88 39.20
Self-Eval Confidence 63.09 30.62 67.22 42.06 73.43 52.01 64.13 27.12
Entropy 33.84 15.16 41.54 19.32 38.31 28.72 56.71 16.15
Semantic Entropy 64.04 21.82 70.89 38.21 66.43 27.90 66.60 26.02
Cluster Entropy 59.86 18.83 65.23 30.25 66.09 30.40 66.99 27.03
MSF 85.66 74.43 86.43 74.42 86.62 73.67 85.10 71.47
LARS 80.50 64.13 85.29 73.09 86.34 72.14 81.37 62.94
HARMONY [Ours] 87.26 76.83 88.71 79.68 86.63 73.75 87.47 76.89
(+1.60) (+2.40) (+2.28) (+5.26) (+0.01) (+0.08) (+2.37) (+5.42)
@@
Baselines We include a range of black-box methods, including length-normalized confidence malinin2020uncertainty, first-token confidence zhao2024first, and self-eval srinivasan2024selective, Entropy, Semantic Entropy, and Cluster Entropy farquhar2024detecting. Semantic Entropy, which measures consistency among semantically similar answers and can be regarded as an alternative implementation of khan2024consistency. Finally, we consider supervised training-based methods, including MSF (Multimodal Selection Function) whitehead2022reliable, which trains an MLP on hidden representations of the prompt and generated answer, and LARS yaldiz2024not, which trains a transformer architecture on the token probabilities predicted by the base model. Details of black-box methods can be found in section 2 and white-box methods in Appendix A.1.
@@ 5.2 Results
5.2.1 UE Performance
LLaVa - 7b LLaVa - 13b Instruct-BLIP Qwen-VL
UE Method ER(%) C@R=10% C@R=20% ER(%) C@R=10% C@R=20% ER(%) C@R=10% C@R=20% ER(%) C@R=10% C@R=20%
@@ 5.2.1 UE Performance
MSF 49.17 43.75 80.17 53.19 50.56 89.52 38.15 32.66 60.08 53.62 46.73 87.77
LARS 49.78 50.65 82.53 53.71 61.83 90.48 37.73 31.27 61.83 53.72 43.14 90.66
HARMONY [Ours] 52.31 60.61 85.59 55.90 64.80 92.23 38.25 36.77 63.32 55.63 59.65 90.74
@@
MSF 21.30 15.00 34.15 23.47 11.90 37.76 13.01 9.15 19.84 16.76 6.01 22.57
LARS 15.72 4.91 21.53 19.14 8.86 29.08 12.82 6.85 17.13 13.09 1.05 15.34
HARMONY [Ours] 21.93 16.37 36.03 24.69 20.24 41.11 13.38 9.23 19.94 18.08 9.64 28.27
We present the comparison results of HARMONY against existing UE baselines on the A-OKVQA and VizWiz datasets in Table 1. We observe that learnable functions such as MSF, LARS, and HARMONY generally perform better across considered datasets and models. Although these methods require a one-time function training cost, they are often more efficient than some of the black-box approaches (cluster entropy, entropy, semantic entropy), which typically require multiple forward passes of the underlying 7B or 13B models. In contrast, HARMONY consists of only 113M trainable parameters and requires a single forward pass to compute the reliability score. Between MSF and LARS, we find that on the VizWiz dataset which contains visually unanswerable questions, MSF surpasses LARS, highlighting the importance of leveraging hidden states information as a multimodal alignment signal in visually uncertain settings. However, for InstructBLIP model with VizWiz dataset, we find that all considered white-box method converge to similar performance levels. Besides that, HARMONY consistently outperforms both LARS and MSF, achieving up to a 5% improvement in AUROC and a 9% improvement in PRR across rest of the models and datasets.
@@
Medical imaging Dataset: We also report the UE performance on medical imaging domain dataset, PathVQA. For this dataset, we find that the output probability based scoring functions perform significantly lower than the learnable functions as shown in 9. Further, HARMONY achieves state-of-the-art performance indicating its effective performance in medical domain as well.
@@ 5.2.2 Selective Prediction Performance
Here, we present the comparison of the learnable scoring functions baselines on the selective prediction task. Before we compare these methods, it is important to mention that for all trainable functions, we select the best model checkpoints based on the AUROC scores. However, a user may choose a different criterion such as ER to achieve higher performance on this downstream task. The objective here is to present a practical use case, and compare the performance of these methods.
Figure 3: An illustration of selective prediction decisions on the A-OKVQA dataset with LLaVa-7b model. In the left-most example, the model generates an incorrect answer, yet both LARS yaldiz2024not MSF whitehead2022reliable choose to answer/predict based on their respective calibration thresholds. In the second example, LARS opts to answer, while MSF correctly abstains. In the right-most example, both methods abstain, whereas HARMONY makes the right prediction for each of these examples.
First, we report how many questions are covered by our method at 10% and 20% risk levels. The larger the number, the better performance. We observe that our proposed method consistently achieves similar or higher coverage across various models, and datasets. We also report ER metric, which represents a better tradeoff between coverage and risk due to a penalty on the incorrectly covered question. For the comparison, we select a threshold for each method giving best ER on the validation split of calibration set, and compute effective reliability using that threshold on the test set. For this metric as well, our method either performs similar or outperforms other methods achieving up to 3.14% higher score. This highlights its potential to yield higher coverage while inuring lower risks. As an example, we present some sample questions in Figure 3 and compare the decision predictions on the trainable functions. While training on either output distributions or hidden representations alone can lead to contradictory or consistently incorrect decisions, leveraging both simultaneously results in better decision functions.
@@
5.2.3 Out-of-Distribution Generalization
To evaluate out-of-distribution (OOD) generalization, we test the LLaVA-13B model trained on the A-OKVQA dataset, which focuses on visual reasoning, by introducing OOD samples from the OKVQA dataset marino2019ok. The experimental setup, illustrated in Figure 4, progressively increases the proportion of OOD samples in the evaluation set. Specifically, the x-axis denotes the percentage of OOD data, where, for example, 33% corresponds to a test mixture containing 33% OKVQA (OOD) samples and 66% A-OKVQA (in-distribution) samples. As the proportion of OOD data increases, we observe a consistent decline in AUROC scores across all supervised baselines, including model confidence based scoring such as Length Normalized Confidence indicating a degradation in uncertainty calibration under distributional shift. However, HARMONY comparatively achieves higher AUROC values, demonstrating relatively better generalization to unseen data distributions.
@@ 5.2.4 Ablation across model layers and input signals
Our hypothesis is that internal layers have a signal of multimodal reliability, however, it is not clear which layer would provide the best signal. Previous works have highlighted that inner layers (layers closer to the input) focus more on extracting lower-level information from the input, while outer layers (layers closer to the model output) are mostly focused on the next token generation azaria2023internal. Therefore, for all the models, we ablate over every fourth layer for both MSF, and our method. We report the best performing layer results in Table 1 for LLaVa-7b. We observe that for LLaVa-7b and 13b models, and Qwen-VL, inner layers (layer 16 and layer 22) yield the best AUROC performance for our method across both datasets. Further, for InstructBLIP, we find the outer-most layer performs the best.
@@
We also conduct an ablation study on partial input signals, generated tokens (Text), token-level probabilities (Prob), and hidden states (HS), using the LLaVA-13B model on the A-OKVQA dataset. We use VisualBERT model as the scoring function transformer architecture for finetuning for the sake of consistency. As shown in 5, we find that text (question and generated answer) without the token probabilities and hidden states yield no significant UE estimate. However, addition of token probabilities accompanying each token text helps the scoring function learn a better UE estimate. Likewise, if we use text and hidden states of the generated tokens as input signals, they yield better results than the MLP based MSF function. Further, combining all three sources of information as proposed in our work yield significantly better reliability estimates.
Figure 5: Effect of calibration data size on the UE performance.
5.2.5 Effect of Calibration Data
Next, we conduct an ablation study to investigate how the function’s performance scales with the amount of calibration data. For this, we vary the calibration set size with LLaVA-13b model on AOKVQA dataset and report AUROC. As shown in Figure 5, increasing the calibration data size consistently enhances the AUROC scores across all learnable baselines. We also observe that HARMONY requires at least 2,000 samples to achieve performance comparable to LNC, and approximately 6,000 samples to surpass existing SOTA baselines. This trend highlights that while learnable approaches benefit significantly from more calibration data, their relative advantage becomes more apparent only once sufficient data are available to capture the diversity of uncertainty patterns in multimodal settings.
Limitations
HARMONY requires labeled calibration data to train the scoring function. While it is needed for other learnable scoring functions (LARS, MSF) as well, the black-box method may not have this limitation. Further, it relies on access to the model’s internal representations, which makes it incompatible with closed-source models. HARMONY may also require architecture-specific adjustments such as for the linear layer projection component, as internal hidden state dimensions vary across VLM families. Addressing these constraints is an important direction for future research on developing reliable and generalizable VLM UE methods.
@@ Conclusion
This work introduces a novel UE method HARMONY for VLMs that effectively combines hidden activation representations with output token probabilities and generated token text. By jointly leveraging model internal states and output beliefs at token level, our proposed framework provides a more holistic reliability assessment, complementing probability-based and representation-based approaches. Our extensive experiments on AOKVQA and VizWiz datasets demonstrate that our method consistently matches or surpasses existing approaches, achieving up to 5% improvement in AUROC and 9% in PRR.
Appendix A Appendix
A.1 Learnable Scoring Function Baselines
In our work, we compare our method, HARMONY, with two existing state-of-the-art baselines, multimodal selection function (MSF) [whitehead2022reliable] and LARS [yaldizdesign]. In the following, we explain these two learnable scoring baselines.
A.1.1 Multimodal Selection Function (MSF)
This work demonstrates the importance of hidden representations in capturing uncertainty in model generation. The key premise of MSF is that using hidden representations of the image, question and answer as input to a simple two-layer MLP can effectively estimate the uncertainty [whitehead2022reliable]. These hidden representations at a given layer have shape [context length,vector dimension] and MSF aggregates them using max or mean pooling across the context length dimension. In autoregressive VLMs, the question tokens usually follow the image, so the hidden representations of the question tokens already encode the relevant visual information through the causal attention mechanism. Furthermore, the large-scale models often contain a substantial number of image tokens (e.g., 576 in LLaVA and 184 in Qwen-VL), many of which are redundant with information already captured in the question-token representations due to image conditioning/prior context. To show this, we performed ablation studies on LLaVA in the AOKVQA and VizWiz datasets and find that the addition of image token hidden representations, reported as a baseline $\text{MSF}^{*}$ in Table 6 does not yield improvement over MSF (without image token hidden representations). To ensure a fair comparison, we, therefore, use the question-token and answer-token representations as input to both the two-layer MLP and our method for all results in Table 1. Under this set of inputs, we show that our method more effectively captures uncertainty by exploiting variation across the context-length dimension, monitoring token-level internal belief signals from hidden representations and output uncertainty from probability embeddings.
Table 6: MSF and $\text{MSF}^{*}$ comparison for LLaVA-7B and LLaVA-13B on AOKVQA and VizWiz datasets.
Dataset UE Method LLaVA-7B LLaVA-13B
AUROC / PRR (%) AUROC / PRR (%)
AOKVQA MSF 78.66 / 67.01 77.64/67.07
$\text{MSF}^{*}$ 78.42 / 66.99 77.61/ 67.01
HARMONY 83.99 / 75.05 83.72 / 77.09
VizWiz MSF 85.66 / 74.43 86.43 / 74.42
$\text{MSF}^{*}$ 85.42 / 73.14 86.15 / 74.38
HARMONY 87.26 / 76.83 88.71 / 79.68
A.1.2 LARS
LARS is a state-of-the-art baseline that captures uncertainty in generation by leveraging token-level uncertainty scores at the output [yaldizdesign]. The method trains a BERT transformer architecture and achieves competitive performance, outperforming various existing black-box approaches. Our work advances this line of research by combining hidden representations with token-level uncertainties, leading to significant performance gains over LARS.
A.2 Training Setup
A.2.1 Dataset Information
We perform evaluations on three datasets: AOKVQA, an image reasoning dataset; VizWiz, a visually answerable or unanswerable image dataset; and PathVQA, a medical imaging dataset. Figure 6 provides example image and question–answer pairs from these datasets to demonstrate the diverse range. For each dataset, we use the provided training split as a calibration dataset. The calibration data is further divided into 80% training and 20% validation partitions to train our scoring function. To evaluate our method, we used the original validation split given with the dataset as the test set. This calibration and test split setup is applied consistently across all three datasets. Details of the resulting train/validation/test splits are provided in Table 7.
Q: How many people can ride this motorcycle at a time?
A: Two
(a) AOKVQA sample
Q: What can is this?
A: Unanswerable
(b) VizWiz sample
Q: Where are liver stem cells (oval cells) located?
A: In the canals of Hering
(c) PathVQA sample
Figure 6: Example question and answer pairs from the three datasets used in our experiments: AOKVQA (subfigure 6(a)), VizWiz (subfigure 6(b)), and PathVQA (subfigure 6(c)).
Table 7: Calibration and Test Data Sizes
Dataset Name Calibration Data (Train + Val) Test
AOKVQA 17056(13645 + 3411) 1145
PathVQA 19653(15723 + 3930) 6259
VizWiz 20523(16418 + 4105) 4319
A.2.2 Model Information
We evaluated our method on four visual language models in our experiments; details of their layers, hidden sizes, and logit sizes are provided in Table 8, giving a brief overview of the sizes of the input vectors used to train the learnable scoring functions. Further, for inference, we use the same prompt for all models: <image >question, please provide a single word or short sentence answer. ASSISTANT:
Table 8: Architecture details of the pretrained visual language models used in our experiments.
Model Name Number of Layers Hidden Dim. Size Vocabulary Size
LLaVa-7B 33 4096 32,000
LLaVa-13B 41 5120 32,000
Instruct-BLIP 25 2048 32,128
Qwen-VL 33 4096 151,936
A.2.3 Hyper-Parameters